OpenStudy (anonymous):

a window glass is 4 ft. 2in by 2ft. 10 in. find its area..

OpenStudy (goformit100):



OpenStudy (anonymous):

then ? length times width ?

OpenStudy (bahrom7893):

convert everything into inches: l=4*12+2 w=2*12+10 A = l*w = (4*12+2)*(2*12+10) inches

OpenStudy (anonymous):

nice .. thanks